Suppose that the price is $40, unit variable cost is $32/unit, and total fixed costs are $3,600.

Required:

a) Compute the unit contribution margin and contribution margin ratio:
unit CM=
CM ratio=  (enter CMR as a fraction of 1, not as %)

b) Write down the CVP relation using CMR: profit as a function of sales revenue.
(fill in the missing numbers in an equation like Profit = 0.35 * Revenue - 50).
Profit =  * Revenue -

c) Based on the CVP relation in (b), what is the breakeven revenue (if necessary, round up to the nearest integer)?


d) Based on the CVP relation in (b), what is the sales revenue required to achieve target profit of $5,000?
Answer #1

a)unit contribution=Sale price - variable expenses

=40 - 32

=8 per unit

CM per unit= Unit Contribution/Sale price

=8/40

=1/5 or .20

b)Profit = 0.20*revenue -3,600

c)on breakeven profit = 0

0=.20*revenue-3,600 or 3,600=0.20*revenue

breakeven revenue = 3,600/0.20

=18,000

d)5,000=0.20*revenue - 3,600

or 5,000+3,600=0.20*revenue

8,600=0.20*revenue

targeted revenue= 8,600/0.20

=43,000